Light is traveling through the two glass slabs as shown below.  

 

 

  •  850 nm 
Nglass1 = 1.460 
Nglass2 = 1.430 

 

 (a) What could be the minimum incident angle if total internal reflection has to occur?      

 

 

 



1
Expert's answer
2021-04-30T11:02:14-0400

"\frac{\sin \left(r\right)}{\sin \left(i\right)}=\frac{n1}{n2}\:"

"Refractive\:index\left(n1\right)\:of\:glass\:1=1.46\theta \:"

"Refractive\:index\left(n2\right)of\:glass\:2=1.43\theta \:"

"\frac{\sin \left(90\right)}{\sin \left(\theta \right)}=\frac{1.46\theta }{1.43\theta }\:"

"\sin \left(\theta \right)=\frac{\sin \left(90\right)1.43\theta }{1.46\theta }"

"\theta =78.4^{\circ }"